Please help us. My daughter and I just cannot figure this out.
Dean has a small garden with a perimeter of 20 ft. and an area of 25 square ft. He wants to triple the dimensions of his garden next year. what will the new perimeter and new area of his larger garden be? explain the changes in perimeter and area.

Answer by mananth(16946)

perimeter of present garden = 20 ft. Area = 25 sq.ft
2(l+w)=20
l+w=10
l*w = 25
l=25/w
substitute l in first equation
25/w+ w= 10
LCD = w
multiply by w
25*w /w +w^2=10w
w^2-10w+25w=0
(w-5)^2=0
w= +/-5
so w=5
so it is a square.
increase size three times
length & width = 15 ft
perimeter = 60 ft
area = 15*15 = 225 sq.ft